# Enhancing Data-Limited Graph Neural Networks by Actively Distilling Knowledge from Large Language Models

### v1

@@
Abstract.
Graphs have emerged as critical data structures for content analysis in various domains, such as social network analysis, bioinformatics, and recommendation systems. Node classification, a fundamental task in this context, is typically tackled using graph neural networks (GNNs). Unfortunately, conventional GNNs still face challenges in scenarios with few labeled nodes, despite the prevalence of few-shot node classification tasks in real-world applications. To address this challenge, various approaches have been proposed, including graph meta-learning, transfer learning, and methods based on Large Language Models (LLMs). However, traditional meta-learning and transfer learning methods often require prior knowledge from base classes or fail to exploit the potential advantages of unlabeled nodes. Meanwhile, LLM-based methods may overlook the zero-shot capabilities of LLMs and rely heavily on the quality of generated contexts. In this paper, we propose a novel approach that integrates LLMs and GNNs, leveraging the zero-shot inference and reasoning capabilities of LLMs and employing a Graph-LLM-based active learning paradigm to enhance GNNs’ performance. Extensive experiments demonstrate the effectiveness of our model in improving node classification accuracy with considerably limited labeled data, surpassing state-of-the-art baselines by significant margins.
@@ Abstract.
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Graphs have become increasingly recognized as one of the powerful data structures to perform real-world content analysis (Chen et al., 2019; Wu et al., 2022; Mohamed et al., 2020; Réau et al., 2023). They are adept at representing complex relationships and uncovering hidden information between objects across various domains. Among various tasks on graphs, node classification stands out as a classic task with broad applications, such as sentiment analysis (Li et al., 2021a) and user attribute inference (Li et al., 2022). Recently, graph neural networks (Kipf and Welling, 2016; Abu-El-Haija et al., 2019; Hamilton et al., 2017), inspired by the idea of convolutional neural networks (CNNs) (Li et al., 2021b), have shown great power in node classification. Generally, GNNs aggregate neighborhood information similar to CNNs, but they exploit the graph’s connectivity structure, facilitating the implicit propagation of information from labeled nodes to unlabeled ones. This strategy has substantially enhanced performance across various benchmark datasets (Wu et al., 2020).
@@
Despite the great success of GNNs in node presentation learning and node classification, they often struggle to generalize effectively when labeled data is scarce. However, in many real-world applications, due to various issues such as labeling cost and privacy issues, one often needs to train a GNN classifier with sparse labels, which is known as few-shot node classification. For example, labeling a large number of web documents can be both costly and time-consuming (Wei et al., 2021; Han et al., 2018); similarly, in social networks, privacy concerns limit access to personal information, leading to a scarcity of attribute labels (Li et al., 2022). Consequently, when confronted with such datasets, GNNs may exhibit poor generalization to unlabeled nodes. To tackle the few-shot learning problem, various methods have been proposed, such as meta-learning (Finn et al., 2017; Ding et al., 2021, 2022), transfer learning (Wan et al., 2021; Zhu et al., 2021), and adversarial reprogramming (Chen et al., 2022). However, they still require a substantial amount of labeled nodes in each class to achieve satisfactory results (Yao et al., 2021) or require auxiliary labeled data to provide supervision.
@@
Recently, Large Language Models (LLMs) have demonstrated their outstanding generalizability in zero-shot learning and reasoning (Chen et al., 2023; Kojima et al., 2022). Several efforts have been taken in introducing LLMs to graph learning, such as pre-processing of textual node attributes or taking textual descriptions of rationales as inputs (Liu et al., 2023; He et al., 2023), leveraging LLMs to construct graph structure (Guo et al., 2024; Jiang et al., 2023), and generating new nodes (Yu et al., 2023). For example, Chen et al. (Chen et al., 2023) first leveraged LLMs as annotators to provide more supervision for graph learning. Yu et al. (Yu et al., 2023) leveraged the generative capacity of LLMs to address the few-shot node classification problem. These works demonstrate that LLMs can enhance GNNs from different perspectives. However, they typically treat LLMS merely as annotators or generators for node classification tasks, overlooking their untapped potentials, such as the capacity to uncover hidden insights within the results and their zero-shot reasoning ability, which could significantly enhance GNNs’ performance for few-shot learning tasks.
@@
In this paper, we introduce a novel few-shot node classification model that enhances GNNs’ capabilities by actively distilling knowledge from LLMs. Unlike previous approaches, our model uses LLMs as “teachers”, capitalizing on their zero-shot inference and reasoning capabilities to bolster the performance of GNNs in few-shot learning scenarios. However, there are two primary challenges: (i) LLMs cannot consistently deliver accurate predictions for all nodes. How to select nodes that LLMs can provide high-quality labels that can benefit GNN most; and (ii) How to effectively distill the knowledge from LLMs to GNNs. To address these challenges, we propose an active learning-based knowledge distillation strategy that selects valuable nodes for LLMs and bridges the gap between LLMs and GNNs. This approach significantly enhances the efficacy of GNNs when labeled data is scarce. We first explore the metrics that affect the correctness of LLMs’ predictions. Then, We employ LLMs as a teacher model and leverage them to perform on the limited training data, generating soft labels for training nodes along with logits and rationales. These outputs are used to supervise GNNs in learning from two perspectives: probability distribution and feature enhancement at the embedding level. In this way, GNNs can learn the hidden information from unlabeled nodes and the detailed explanation provided by LLMs. Furthermore, we introduce a novel Graph-LLM-based active learning approach to establish a connection between LLMs and GNNs, which effectively select nodes for which GNNs fail to provide accurate pseudo-labels but LLMs can offer reliable pseudo-labels, thereby enabling GNNs to leverage the zero-shot capabilities of LLMs and enhance their performance with limited data. Afterward, the selected pseudo-labels are merged with the true labels to train the final few-shot node classification model. In summary, the major contributions of our paper are:
  • We analyze the correctness of LLMs’ predictions and conduct preliminary experiments to explore metrics that affect the correctness of LLMs’ predictions.
@@
  • We design and implement a Graph-LLM-based active learning paradigm to enhance the performance of GNNs. This is achieved by identifying nodes for which GNNs struggle to generate reliable pseudo labels, yet LLMs can provide dependable predictions, which leverage the zero-shot learning and reasoning ability of LLMs to enhance the performance of GNNs.
@@
  • Extensive experiments on various benchmark datasets demonstrate the effectiveness of the proposed framework for node classification tasks even with limited labeled data.
2. Related Work
In this section, we introduce related works, including graph neural networks, few-shot node classification, LLMs for text-attributed graphs, and active learning.
2.1. Graph Neural Networks
@@
Graph Neural Networks (GNNs) have garnered widespread attention for their effective exploitation of graph structure information. There are two primary types of GNNs: spectral-based and spatial-based. Kipf and Welling (Kipf and Welling, 2016) followed the idea of CNNs and proposed the Graph Convolutional Network (GCN) to aggregate information within the spectral domain using graph convolution. Different from GCN, Graph Attention Network (GAT) (Veličković et al., 2018) and GraphSAGE (Hamilton et al., 2017) emerged as spatial-based approaches. GAT applies the attention mechanism to learn the importance of the neighbors when aggregating information. GraphSAGE randomly samples the number of neighbors of a node and aggregates information from these local neighborhoods. Despite their extensive application across various domains, GNNs often face challenges due to limited labeled data. Existing convolutional filters or aggregation mechanisms struggle to effectively propagate labels throughout the entire graph when only few labeled data points are available (Li et al., 2018).
2.2. Few-shot Node Classification
@@
In real-world graph learning tasks, obtaining high-quality labeled samples can be particularly challenging due to various factors such as the high cost involved in annotation processes or the limited access to node information. Thus, researchers proposed different methods to improve the performance of GNNs with only few labeled data. Most recent advancements in few-shot node classification (FSNC) models have mainly developed from two approaches: metric-learning and meta-learning. Metric-learning models aim to learn a task-invariant metric applicable across all tasks to facilitate FSNC (Hao et al., 2019; Jiang et al., 2020). Prototypical network (Snell et al., 2017) and relation network (Sung et al., 2018) are two classic examples, where the former uses the mean vector of the support set as a prototype and calculates the distance metrics to classify query instances, and the latter trains a neural network to learn the distance metric between the query and support set. Meta-learning models use task distributions to conduct meta-training, learning shared initialization parameters that are then adapted to new tasks during meta-testing (Ding et al., 2021; Zhou et al., 2019; Huang and Zitnik, 2020). These approaches have demonstrated effectiveness compared to metric-learning, which often struggles due to task divergence issues. However, meta-learning requires significant amounts of data for meta-training, sourced from the same domain as meta-testing, thereby severely limiting its practical applicability. Different from metric-learning and meta-learning models, in this paper, we propose to distill knowledge from LLMs to GNNs, leverage the LLMs’ zero-shot ability and reasoning ability to improve GNNs for few-shot node classification.
2.3. LLMs for Text-Attributed Graphs
Recently, Large language models (LLMs) have garnered widespread attention and experienced rapid development, emerging as a hot topic in the artificial intelligence area. By training on extensive datasets, these models have learned to understand and generate natural language, demonstrating remarkable capabilities across a variety of tasks including text generation (Li et al., 2024b), question answering (Li et al., 2024a), translation (Yao et al., 2023), as well as more complex reasoning and creative tasks. Researchers are integrating Large Language Models (LLMs) into diverse fields such as computer vision, time series analysis, and Graph Neural Networks (GNNs), expanding their applicability beyond natural language processing.
@@ 2.3. LLMs for Text-Attributed Graphs
Within the graph domain, LLMs show their generalizability in dealing with Text-Attributed Graphs (TAGs). Chen et al. (Chen et al., 2023) demonstrated the power of LLMs’ zero-shot ability on node classification tasks. Moreover, LLMs also demonstrate their power in providing rationales to enhance node features (He et al., 2023) and construct edges in graphs (Sun et al., 2023). Liu et al. (Liu et al., 2023) further proposed OFA to encode all graph data into text and leverage LLMs to make predictions on different tasks. Despite their remarkable proficiency in understanding text, LLMs still face limitations when it comes to processing graph-structured data. Therefore, leveraging LLMs’ zero-shot ability and integrating them with GNNs has emerged as the latest state-of-the-art approach in text-attributed graph learning (Chen et al., 2023).
2.4. Active Learning
@@
Active learning (AL) (Aggarwal et al., 2014; Shen et al., 2017; Bachman et al., 2017; Cai et al., 2017; Gao et al., 2018; Wu et al., 2019) is a widely adopted approach across various domains for addressing the issue of label sparsity. The core concept involves selecting the most informative instances from the pool of unlabeled data. Recently, many works (Cai et al., 2017; Gao et al., 2018; Hu et al., 2020) integrate GNNs with AL to improve the representative power of graph embeddings. However, how to leverage AL to build connections between LLMs and GNNs and improve the performance of GNNs has emerged as a problem. Chen et al. (Chen et al., 2023) first leverage active learning to select nodes that are close to the cluster center under the label-free setting and use LLM as an annotator to create labels for these nodes. However, their approach simply leverages LLMs to annotate nodes and ignores the benefits of unlabeled nodes and the zero-shot reasoning ability of LLMs. And, under the few-shot setting, GNNs themselves can provide relatively high-quality pseudo-labels for those nodes close to the cluster center, which waste resources if we use LLMs to generate pseudo-labels for those nodes. Moreover, prior research primarily concentrated on selecting data with the highest confidence score during the AL process. In our work, instead of focusing on nodes where GNNs have high confidence, we prioritize nodes where GNNs struggle to provide pseudo-labels with high confidence scores but LLMs can provide reliable predictions. This approach is motivated by our integration of LLMs as teacher models to enhance the performance of GNNs by leveraging LLMs’ zero-shot pseudo-labeling and reasoning ability. Through active learning, we integrate LLMs into GNNs, enabling LLMs to instruct GNNs with data that GNNs find challenging to label confidently.
3. Preliminaries
@@
3.1. Notations
We use $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to denote a graph, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is a set of $N$ nodes and $\mathcal{E}$ is a set of edges. We use $\mathbf{A}$ to denote the adjacency matrix, where $\mathbf{A}_{ij}=1$ means nodes $v_{i}$ and $v_{j}$ are connected; otherwise $\mathbf{A}_{ij}=0$. The text-attributed graph can be defined as $\mathcal{G}_{T}=(\mathcal{V},\mathbf{A},\mathcal{X})$, where $\mathcal{X}=\{\mathcal{X}_{1},\mathcal{X}_{2},\cdots,\mathcal{X}_{N}\}$ denotes the set of raw texts and can be encoded as text embeddings $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{N}\}$. In semi-supervised learning, The node set $V$ can be divided into two different sets: (1) the labeled node set $\mathcal{V}_{l}$ and (2) the unlabeled node set $\mathcal{V}_{u}$. Moreover, we use $\mathcal{V}_{S}$ to denote the labeled node set including both original labeled data and the data selected through active learning, and use $\mathcal{Y}$ to represent the label set, where $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{N}\}$.
Figure 1. Preliminary experiments: the impact of the degrees (left) and homophily ratio (right) to LLMs
3.2. Understanding LLM’s Capability
As the sparse label challenges GNN, in this paper, we aim to imbue GNNs with the zero-shot learning prowess of LLMs, thereby elevating their performance in scenarios with limited labeled data. However, LLMs might be good at classifying certain nodes while performing poorly on other nodes. Thus, it is important to identify nodes that LLMs can provide superior pseudo-labels with rationales, whereas GNNs cannot, which can better enhance GNNs’ performance. Hence, we first conduct preliminary experiments to understand the key metrics that are pivotal for LLMs in generating reliable pseudo-labels.
LLMs may benefit from various metrics to perform node classification well. Particularly, in graph $\mathcal{G}$, certain metrics exert a more pronounced influence on the correctness of LLM predictions on nodes, which include: 1) node features, such as the title and abstract in the citation network; 2) degree of a node, and 3) the homophily ratio. As experimentally demonstrated in (Huang et al., 2023; He et al., 2023), LLMs can provide better classification results with nodes that contain richer information. Therefore, we decided to leverage LLMs as a teacher model to provide enhanced rationales to teach GNNs from the feature perspective at the embedding level. However, when using active learning to identify valuable nodes, we primarily rely on the original node features. Both degrees and the homophily ratio are important for a node. The former indicates how many nodes will be affected by a node, while the latter suggests that the node tends to connect with others having similar features. Therefore, it’s crucial to examine how the degree and the homophily ratio affect the performance of LLMs’ predictions.

### v2

@@
Abstract
Graphs are pervasive in the real-world, such as social network analysis, bioinformatics, and knowledge graphs. Graph neural networks (GNNs) have great ability in node classification, a fundamental task on graphs. Unfortunately, conventional GNNs still face challenges in scenarios with few labeled nodes, despite the prevalence of few-shot node classification tasks in real-world applications. To address this challenge, various approaches have been proposed, including graph meta-learning, transfer learning, and methods based on Large Language Models (LLMs). However, traditional meta-learning and transfer learning methods often require prior knowledge from base classes or fail to exploit the potential advantages of unlabeled nodes. Meanwhile, LLM-based methods may overlook the zero-shot capabilities of LLMs and rely heavily on the quality of generated contexts. In this paper, we propose a novel approach that integrates LLMs and GNNs, leveraging the zero-shot inference and reasoning capabilities of LLMs and employing a Graph-LLM-based active learning paradigm to enhance GNNs’ performance. Extensive experiments demonstrate the effectiveness of our model in improving node classification accuracy with considerably limited labeled data, surpassing state-of-the-art baselines by significant margins.
Index Terms:
@@ Abstract
I Introduction
@@
Graphs have become increasingly recognized as one of the powerful data structures to perform real-world content analysis [1, 2, 3]. They are adept at representing complex relationships and uncovering hidden information between objects across various domains. Among various tasks on graphs, node classification stands out as a classic task with broad applications, such as sentiment analysis [4] and user attribute inference [5]. Recently, graph neural networks [6, 7, 8] have shown great power in node classification. Generally, GNNs adopt the message-passing mechanism, which updates a node’s representation by aggregating its neighbors’ information, facilitating the implicit propagation of information from labeled nodes to unlabeled ones. This strategy has substantially enhanced performance across various benchmark datasets [9].
@@
Despite the great success of GNNs in node presentation learning and node classification, they often struggle to generalize effectively when labeled data is scarce. However, in many real-world applications, due to various reasons such as labeling cost and privacy issues, one often needs to train a GNN classifier with sparse labels, which is known as few-shot node classification. For example, labeling a large number of web documents can be both costly and time-consuming [10, 11]; similarly, in social networks, privacy concerns limit access to personal information, leading to a scarcity of attribute labels [5]. Consequently, when confronted with such datasets, GNNs may exhibit poor generalization to unlabeled nodes. To tackle the few-shot learning problem, various methods have been proposed, such as meta-learning [12, 13, 14], transfer learning [15, 16], and adversarial reprogramming [17]. However, they still require a substantial amount of labeled nodes in each class to achieve satisfactory results [18] or require auxiliary labeled data to provide supervision.
@@
Recently, Large Language Models (LLMs) have demonstrated their outstanding generalizability in zero-shot learning and reasoning [19, 20]. Several efforts have been taken in introducing LLMs to graph learning, such as pre-processing of textual node attributes or taking textual descriptions of rationales as inputs [21, 22], leveraging LLMs to construct graph structure [23, 24], and generating new nodes [25]. For example, Chen et al. [19] first leveraged LLMs as annotators to provide more supervision for graph learning. Yu et al. [25] leveraged the generative capability of LLMs to address the few-shot node classification problem. These works demonstrate that LLMs can enhance GNNs from different perspectives. However, they typically treat LLMS merely as annotators or generators for node classification tasks, overlooking their untapped potentials, such as the capacity to uncover hidden insights within the results and their zero-shot reasoning ability, which could significantly enhance GNNs’ performance for few-shot learning tasks.
@@
In this paper, we introduce a novel few-shot node classification model that enhances GNNs’ capabilities by actively distilling knowledge from LLMs. Unlike previous approaches, our model uses LLMs as “teachers”, capitalizing on their zero-shot inference and reasoning capabilities to bolster the performance of GNNs in few-shot learning scenarios. However, there are two primary challenges: (i) LLMs cannot consistently deliver accurate predictions for all nodes. How to select nodes that LLMs can provide high-quality labels that can benefit GNN most; and (ii) How to effectively distill the knowledge from LLMs to GNNs. To address these challenges, we propose an active learning-based knowledge distillation strategy that selects valuable nodes for LLMs and bridges the gap between LLMs and GNNs. This approach significantly enhances the efficacy of GNNs when labeled data is scarce. We first explore the metrics that affect the correctness of LLMs’ predictions. Then, we employ LLMs as a teacher model and leverage them to perform on the limited training data, generating soft labels for training nodes along with logits and rationales. These outputs are used to supervise GNNs in learning from two perspectives: probability distribution and feature enhancement at the embedding level. In this way, GNNs can learn the hidden information from unlabeled nodes and the detailed explanation provided by LLMs. Furthermore, we introduce a novel Graph-LLM-based active learning approach to establish a connection between LLMs and GNNs, which effectively select nodes for which GNNs fail to provide accurate pseudo-labels but LLMs can offer reliable pseudo-labels, thereby enabling GNNs to leverage the zero-shot capabilities of LLMs and enhance their performance with limited data. Afterward, the selected pseudo-labels are merged with the true labels to train the final few-shot node classification model. Our major contributions are:
@@
  • We design and implement a Graph-LLM-based active learning paradigm to enhance the performance of GNNs. This is achieved by identifying nodes for which GNNs struggle to generate reliable pseudo labels, yet LLMs can provide dependable predictions, which leverage the zero-shot ability of LLMs to enhance the performance of GNNs.
@@
  • Extensive experiments on various benchmark datasets demonstrate the effectiveness of our proposed framework for node classification tasks with limited labeled data.
II Related Work
@@
Graph Neural Networks Graph Neural Networks (GNNs) have garnered widespread attention for their effective exploitation of graph structure information. There are two primary types of GNNs: spectral-based and spatial-based. Kipf and Welling [6] followed the idea of CNNs and proposed the Graph Convolutional Network (GCN) to aggregate information within the spectral domain using graph convolution. Different from GCN, Graph Attention Network (GAT) [26] and GraphSAGE [8] emerged as spatial-based approaches. GAT applies the attention mechanism to learn the importance of the neighbors when aggregating information. GraphSAGE randomly samples the number of neighbors of a node and aggregates information from these local neighborhoods. Despite their extensive application across various domains, GNNs often face challenges due to limited labeled data. Existing convolutional filters or aggregation mechanisms struggle to effectively propagate labels throughout the entire graph when only few labeled data points are available [27].
@@
Few-shot Node Classification In real-world graph learning tasks, obtaining high-quality labeled samples can be particularly challenging due to various factors such as the high cost involved in annotation processes or the limited access to node information. Thus, researchers proposed different methods to improve the performance of GNNs with only few labeled data. Most recent advancements in few-shot node classification (FSNC) models have mainly developed from two approaches: metric-learning and meta-learning. Metric-learning models aim to learn a task-invariant metric applicable across all tasks to facilitate FSNC [28, 29]. Prototypical network [30] and relation network [31] are two classic examples, where the former uses the mean vector of the support set as a prototype and calculates the distance metrics to classify query instances, and the latter trains a neural network to learn the distance metric between the query and support set. Meta-learning models use task distributions to conduct meta-training, learning shared initialization parameters that are then adapted to new tasks during meta-testing [13, 32, 33]. These approaches have demonstrated effectiveness compared to metric-learning, which often struggles due to task divergence issues. However, meta-learning requires significant amounts of data for meta-training, sourced from the same domain as meta-testing, thereby severely limiting its practical applicability. Different from metric-learning and meta-learning models, in this paper, we propose to distill knowledge from LLMs to GNNs, leverage the LLMs’ zero-shot ability and reasoning ability to improve GNNs for few-shot node classification.
@@
LLMs for Text-Attributed Graphs Recently, LLMs have garnered widespread attention and experienced rapid development, emerging as a hot topic in the artificial intelligence area. Within the graph domain, LLMs show their generalizability in dealing with Text-Attributed Graphs (TAGs). Chen et al. [19] demonstrated the power of LLMs’ zero-shot ability on node classification tasks. Moreover, LLMs also demonstrate their power in providing rationales to enhance node features [22] and construct edges in graphs [34]. Liu et al. [21] further proposed OFA to encode all graph data into text and leverage LLMs to make predictions on different tasks. Despite their remarkable proficiency in understanding text, LLMs still face limitations when it comes to processing graph-structured data. Therefore, leveraging LLMs’ zero-shot ability and integrating them with GNNs has emerged as the latest state-of-the-art approach in text-attributed graph learning [19].
@@
Active Learning Active learning (AL) [35, 36, 37, 38, 39] is a widely adopted approach across various domains for addressing the issue of label sparsity. The core concept involves selecting the most informative instances from the pool of unlabeled data. Recently, many works [37, 38, 40] integrate GNNs with AL to improve the representative power of graph embeddings. However, how to leverage AL to build connections between LLMs and GNNs and improve the performance of GNNs has emerged as a problem. Chen et al. [19] first leverage active learning to select nodes that are close to the cluster center under the label-free setting and use LLM as an annotator to create labels for these nodes. However, their approach simply leverages LLMs to annotate nodes and ignores the benefits of unlabeled nodes and the zero-shot reasoning ability of LLMs. And, under the few-shot setting, GNNs themselves can provide relatively high-quality pseudo-labels for those nodes close to the cluster center, which waste resources if we use LLMs to generate pseudo-labels for those nodes. Moreover, prior research primarily concentrated on selecting data with the highest confidence score during the AL process. In our work, instead of focusing on nodes where GNNs have high confidence, we prioritize nodes where GNNs struggle to provide pseudo-labels with high confidence scores but LLMs can provide reliable predictions. This approach is motivated by our integration of LLMs as teacher models to enhance the performance of GNNs by leveraging LLMs’ zero-shot pseudo-labeling and reasoning ability. Through active learning, we integrate LLMs into GNNs, enabling LLMs to instruct GNNs with data that GNNs find challenging to label confidently.
III Preliminaries
@@
Notations We use $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to denote a graph, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is a set of $N$ nodes and $\mathcal{E}$ is a set of edges. We use $\mathbf{A}$ to denote the adjacency matrix, where $\mathbf{A}_{ij}=1$ means nodes $v_{i}$ and $v_{j}$ are connected; otherwise $\mathbf{A}_{ij}=0$. The text-attributed graph can be defined as $\mathcal{G}_{T}=(\mathcal{V},\mathbf{A},\mathcal{X})$, where $\mathcal{X}=\{\mathcal{X}_{1},\mathcal{X}_{2},\cdots,\mathcal{X}_{N}\}$ denotes the set of raw texts and can be encoded as text embeddings $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{N}\}$. In semi-supervised learning, The node set $\mathcal{V}$ can be divided into two different sets: (1) the labeled node set $\mathcal{V}_{l}$ and (2) the unlabeled node set $\mathcal{V}_{u}$. Moreover, we use $\mathcal{V}_{S}$ to denote the labeled node set including both original labeled data and the data selected through active learning, and use $\mathcal{Y}$ to represent the label set, where $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{N}\}$.
Figure 1: Preliminary experiments: the impact of the degrees (left) and homophily ratio (right) to LLMs
III-A Understanding LLM’s Capability
As the sparse label challenges GNN, in this paper, we aim to imbue GNNs with the zero-shot learning prowess of LLMs, thereby elevating their performance in scenarios with limited labeled data. However, LLMs might be good at classifying certain nodes while performing poorly on other nodes. Thus, it is important to identify nodes that LLMs can provide superior pseudo-labels with rationales, whereas GNNs cannot, which can better enhance GNNs’ performance. Hence, we first conduct preliminary experiments to understand key factors pivotal for LLMs in generating reliable pseudo-labels.
LLMs may benefit from various metrics to perform node classification well. Particularly, in graph $\mathcal{G}$, certain metrics exert a more pronounced influence on the correctness of LLM predictions on nodes, which include: 1) the degree of a node, and 2) the homophily ratio. Both degrees and the homophily ratio are important for a node. The former indicates how many nodes will be affected by a node, while the latter suggests that the node tends to connect with others having similar features. Therefore, it’s crucial to examine how the degree and the homophily ratio affect the performance of LLMs’ predictions.